OpenStudy (sleepyjess):

Need help on 2 exponential growth problems please :)

OpenStudy (sleepyjess):

The population of rabbits in a forest follows the law of exponential growth. There are 150 rabbits present initially and there are 300 after 3 years. (a) Write a function that represents the number of rabbits after t years. (b) How many rabbits will there be in 10 years? How do I get the function?

OpenStudy (aveline):

\[y=a(1+r)^{x}\]

OpenStudy (irishboy123):

start with the idea that your equation is in the form \(N = N_o e^{kt}\)

OpenStudy (aveline):

That's your exponential growth equation. a is the initial amount, r is the growth rate, and x is the number of time intervals that has passed

OpenStudy (sleepyjess):

Okay, I remember that equation from the lesson @IrishBoy123 , if I remember right, with the information given, it would become 300 = 150\(e^{3k}\)?

OpenStudy (sleepyjess):

Then I need to get \(e^{3k}\) on the right by itself?

OpenStudy (irishboy123):

yes, just put t = 0 and t = 3 in with your other conditions and solve them so, for t = 0 you have \(150 = N_o e^{0} = N_o\) for t = 3 you have \(300 = 150 e^{3k}\). solve that using natural logs but @Aveline 's idea looks good.

OpenStudy (sleepyjess):

Just to make sure I remember how to solve with natural logs, I would take ln of both sides, ln 2 = 3k ln e, ln e cancels out leaving ln 2 = 3k?

OpenStudy (irishboy123):

or \(\dfrac{300}{150} = e^{3k}\) \(\ln \dfrac{300}{150} = 3k\) same as you

OpenStudy (sleepyjess):

Okay, thank you! :)

OpenStudy (sleepyjess):

I think I can do the second problem on my own now :)